I do not know how to solve (0,0) m= 5/11. The answer must be written in Ax+By=C
Answer by stanbon(75887)

I do not know how to solve (0,0) m= 5/11. The answer must be written in Ax+By=C
---
You want the equation of a line that passes thru (0,0)
and has a slope of 5/11.
----
The (0,0) tell you the y-intercept is zero.
---
The Form of the equation is y = mx + b
where m is the slope and b is the intercept.
-----------
Equation:
y = (5/11)x + 0
---
Putting that in the form Ax+By = C
you get:
11y = 5x
5x - 11y = 0
====================
Cheers,
Stan H.
====================